OpenStudy (0_0youscareme):

evaluate 4(5*1 divided by 20). Name the property used in each step

OpenStudy (0_0youscareme):

@jim_thompson5910

OpenStudy (0_0youscareme):

@mathstudent55

OpenStudy (0_0youscareme):

@ganeshie8

jimthompson5910 (jim_thompson5910):

how far did you get here?

OpenStudy (0_0youscareme):

i know how 2 solve it i just dont know what properties are used

jimthompson5910 (jim_thompson5910):

can you write out how to solve it? Then I'll help out with the properties used

OpenStudy (0_0youscareme):

well i have 2 use pemdas for this

OpenStudy (0_0youscareme):

so it would be 5*1=5 4(5 divided by 20)

OpenStudy (0_0youscareme):

and then 5 divided by 20 witch gives u .25 4(.25) then multiply 4*0.25 witch gives u 1

OpenStudy (0_0youscareme):

right?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

you can also do it like this 4(5/20) would turn into (4*5)/20 then 4*5 = 20 so we really have 20/20 = 1

jimthompson5910 (jim_thompson5910):

as for the properties used, there isn't some really fancy name for it. You're just doing simple multiplication and division. I'm not sure what the teacher is trying to get you to do here I guess you could say "associative property" when rearranging terms basically 4(5/20) = 4*5*(1/20) = (4*5)*(1/20) but that seems like we're making things more complicated than they have to be

OpenStudy (0_0youscareme):

i put multiplication and division the first time and i got it wrong so yea

jimthompson5910 (jim_thompson5910):

Is there a list of choices they want you to pull from?

OpenStudy (0_0youscareme):

no

OpenStudy (0_0youscareme):

but tysm for ur help

jimthompson5910 (jim_thompson5910):

no problem